Fill in the missing information from the balance sheet. Stuart Company Balance Sheet As of December 31, 2017 (amounts in thousands) Cash Liabilities 2,900 Other Assets 3,000 Equity 2,100 Total Assets Total Liabilities & Equity What is the value for Cash? Please specify your answer in the same units as the balance sheet.
Answer #1

All the amounts are in thousands.

Total liabilities and equity = Liabilities + Equity = 2,900 + 2,100 = 5,000

Total assets = Total liabilities and equity

\thereforeTotal assets = 5,000

Cash = Total assets - Other assets = 5,000 - 3,000 = 2,000